A 100 mL sample of an acetic acid/sodium acetate bffer
is diluted with 100 mL of distilled water. What is the effect of the dilution on the pH of the buffer? What is the e ect of the dilution on the buffer capacity?
Answer #1

A buffer is a mixture of a weak acid (in this case the acetic acid) and its conjguate base (the acetate)

it is shown in equations:

HA(aq) <-> H+(aq) + A-(aq)

NaA(aq) --> Na+(aq) + A-(aq)

note that A- is present in both cases so.... this forms a buffer due to the equilibrium of

K = [H+][A-]/[HA]

then

adding acid or base will make a small change of pH, due to the equilibrium shown before.

if we dilute those 100 mL of buffer with another 100 mL of water,

VT = V1+V2 = 100+100 = 200 mL

the new volume is twice.

Concentrations

C = mol/V

are decreased by a factor of two, since V is now 2 times larger, then C is 2 times smaller.

BUT!

the equation that best describes the buffer is that of Henderson Hasselbalch

which states that

pH = pKa + log([A-]/[HA])

since pKa is a constant value for acetic acid

then ignore it

pH is dependent of log([A-]/[HA]) only

and since

[A-] = halved

[HA] = halved

the effect is actually

log([A-]/[HA]) = (1/2 A-) / (1/2 HA) = 1*(A/HA)

that is,

there is no EFFECT in the buffering capacity...